OpenStudy (anonymous):

The total price of 4 equally priced markers is $6.40. If 10 or more markers are purchased, the cost per marker is decreased by .25 for the entire order. How much will 12 of these markers cost?

OpenStudy (anonymous):

Well you know a group of 4 costs $6.40. Multipl that three times to gt (12 markers) which brings the cost to $19.20 before the decrease. Can you figure it now?

OpenStudy (anonymous):

Still a bit confused.

OpenStudy (anonymous):

U said th cost pr markr is decreased by .25 if mor than 10 are bought. So, I believe that comes up to $1.25 which u can - from $19.20 which gives you $17.95

OpenStudy (anonymous):

Hope that helps ;)

OpenStudy (anonymous):

I came up with a different answer. @sakigirl is that the right answer?

OpenStudy (anonymous):

I'm not sure, what's your answer?

OpenStudy (anonymous):

@romero

OpenStudy (anonymous):

well you said that when you bought 4 markers (let x represent price of markers) the total was 6.40. 4*x=6.40 x=6.40/4 x=1.60

OpenStudy (anonymous):

So multiplying 6.40 x 12 right?

OpenStudy (anonymous):

BazBendell's answer is incorrect, I'll tell you that much.

OpenStudy (anonymous):

When you buy 10 or more you can discount .25 You can do this by initial taking .25 from each markers initial price or by the overall price. From initial price: 1.60 *.25 = 0.40 (this is the discount) 1.60- 0.40 = 1.20 (This is the price of each marker when buying 10 or more) This is the price we use when buying 10 or more 1.20 * 12 = ? (12 is how many marker you bought) 12.00 +2.40= 14.40 (didn't use a calculator :/)

OpenStudy (anonymous):

Answer is the same in both cases!! There is always one way to do things in Math! Don't forget that. Second form requires more math but you can see and compare how much it would have been without the discount and how much you save.

OpenStudy (anonymous):

@romero So isn't BazBendell right? The question was never asking for the discount price :P

OpenStudy (anonymous):

Nope that person is not right. My first way of doing it requires less work and goes straight to the answer which is "How much will 12 of these markers cost" Which is 14.40

OpenStudy (anonymous):

14.40 is not in the answer choices @romero

OpenStudy (anonymous):

What are the answer choices?

OpenStudy (anonymous):

@romero 13.50 16.20 18.20 19.20 22.20

OpenStudy (anonymous):

Are you sure he is wrong?

OpenStudy (anonymous):

His answer is not one of the answer choices.

OpenStudy (anonymous):

Oh, oops. I thought he said the answer was 19.20

OpenStudy (anonymous):

Do you know what might have gone wrong?

OpenStudy (anonymous):

No I don't see it. Maybe the answer choices are wrong.

OpenStudy (anonymous):

Highly unlikely.

OpenStudy (anonymous):

But not impossible. Maybe you wrote the problem wrong. With the given information and question you asked I came up with 14.40